(4/x)-(3/(x+1))=7 slove for x
Answer by stanbon(75887)

(4/x)-(3/(x+1))=7 solve for x
----------
Multiply by the lcm of the denominators = x(x+1)
4(x+1)-3x = 7x(x+1)
4x+4-3x = 7x^2+7x
7x^2+6x-4=0
x=[-6+-sqrt(36-4*7*-4)]/14
x=[-6+-sqrt148]/14
x=[-6+-2sqrt37]/14
x=[-3+-sqrt37]/7
===========
Cheers,
Stan H.




==============
Cheers,
Stan H.